OpenStudy (anonymous):

Calculate the surface area of a sphere whose radius is 8mm

OpenStudy (anonymous):

surface area of a sphere 4(3.14)r^2 = 4(3.14)(8)^2 =803.84

OpenStudy (ash2326):

surface area is given by the relation \[ S= 4* \pi * r^2\] \[ S=4 *\ pi* 8^2\] \[ S=803.84 mm^2\]

OpenStudy (anonymous):

thanks!